a solid conducting sphere with an area of 16m2 has total charge of 4"\mu C." Determine the electric field through the Gaussian surface


1
Expert's answer
2022-02-28T09:59:42-0500

The Gauss's law says

"\Phi=\oint EdA=\frac{q}{\epsilon_0}"

"\Phi=\frac{4*10^{-6}}{8.85*10^{-12}}=4.5*10^5\:\rm N\cdot m^2/C"